Part A Pyridine is a weak base that is used in the manufacture of pesticides and...

Part A

Pyridine is a weak base that is used in the manufacture of pesticides and plastic resins. It is also a component of cigarette smoke. Pyridine ionizes in water as follows:

C5H5N+H2O⇌C5H5NH++OH−

The pKb of pyridine is 8.75. What is the pH of a 0.185 M solution of pyridine? (Assume that the temperature is 25 ∘C.)

Part B

Benzoic acid is a weak acid that has antimicrobial properties. Its sodium salt, sodium benzoate, is a preservative found in foods, medications, and personal hygiene products. Benzoic acid ionizes in water:

C6H5COOH⇌C6H5COO− + H+

The pKa of this reaction is 4.2. In a 0.62 M solution of benzoic acid, what percentage of the molecules are ionized?
Answer #1

? (shs Nht + oH A) CS MEN + H2O 0.185M (initial) 0.1854 0.185. 0.18582 laurel (& is very small for weak bad) 0.185 (1-0) Kb =0.620 0.62x B) GH5 ook ons loo + H+ 0.62 m 0 0 (initial) 0.62 (1-0) ka = 10.620) (0.62d) 0.620² 0.6211-«) 1-821 (& is very sm